Curriculum vitae of Imre Csiszár

Imre Csiszár was born in Miskolc, Hungary, on February 7, 1938. He received the diploma in mathematics from the L. Eötvös University, Budapest, and the Doctor of Mathematical Science degree from the Hungarian Academy of Sciences in 1977. In 1990 he was elected Corresponding Member and in 1995 Regular Member of the Hungarian Academy of Sciences.

Since 1961, he has been with the Mathematical Institute of the Hungarian Academy of Sciences, and has been Head of the Information Theory Group there since 1968. He is also Professor of Mathematics at the L. Eötvös University, Budapest and the Technical University, Budapest. He has held Visiting Professorships at various Universities including the University Bielefeld, Germany, the University of Maryland at College Park, MD, Stanford University, Stanford, CA, University of Virginia, Charlottesville, VA, and the University of Tokyo, Japan. His research interests are centered on Shannon theory and on applications of information theory in mathematics, primarily in probability and statistics. He is cooauthor of the book Information Theory: Coding Theorems for Discrete Memoryless Systems (New York: Academic Press, 1981).

Dr. Csiszár is President of the Hungarian Mathematical Society, and Fellow of the IEEE. He has been recepient of the 1988 Paper Award of the IEEE Information Theory Society, the Academy Award for Interdisciplinary Research of the Hungarian Academy of Sciences in 1989, and the Shannon Award of the IEEE Information Theory Society in 1996.